OpenStudy (anonymous):

Explain, complete sentences, the process of using Cramer’s Rule to solve the following system of equations. Provide the solution to the system. 8x + 3y = 14 12x + 4y = 17

OpenStudy (anonymous):

Set up one determinant for D and one for Dy. The values for determinant D are taken from the coefficients of the system (8 and 3 in row 1, 12 and 4 in row 2): D= |8 3| |12 4| Dy is set up the same way, but substitute the constants (14 and 17) for the coefficients of y in the determinant: Dy= |8 14| |12 17| Solve the determinants: D= |a b| |c d| Dy= |a e| |c f| D is solved by (a)(d)-(b)(c); and Dy is solved by (a)(f)-(e)(c) D= (8)(4)-(3)(12) = 32 - 36 = -2 Dy= (8)(17)-(14)(12) = 136 - 168 = -32 y=Dy/D y=-32/-2 = 16

OpenStudy (anonymous):

Thank you, I figured out what I was doing wrong.